# Equivariant Symmetry-Aware Head Pose Estimation for Fetal MRI

### v4

@@ Appendix E Datasets
This dataset consists of whole-uterus EPI volumes in 153 pregnant patients (3T, Siemens, 3mm isotropic resolution, TR=2.9-4s, TE=32-47 ms, flip angle=90). Fetus GA ranges from 18 to 38 weeks, with mean GA $29.0\pm 5.0$ weeks. The average FOV size is $300\!\times\!300\!\times\!156$$m$.
@@
This dataset consists of fetal brain fMRI scans in 245 pregnant patients (3T, Philips Achieva, 2.2mm isotropic resolution, TR=2.2s, TE=60ms, flip angle=90) [karolis2025]. Fetus GA ranges from 20 to 38 weeks, with mean GA $28.9\pm 3.8$ weeks. The average FOV size is $143\!\times\!151\!\times\!108$mm. Volumes are cropped around the fetal head and are not whole-uterus volumes. Our manually annotated brain/eyes segmentations and poses for the dHCP dataset are publicly available on the project website for full transparency and future use.
@@
This dataset consists of whole-uterus EPI volumes in 60 pregnant patients (3T, Siemens, 1.76-3.5mm voxels, TR=2.2-5s, TE=37ms, flip angle=90). There are 10 subjects for each GA in the range of 18 to 23 weeks. The average FOV size is $301\!\times\!301\!\times\!129$$m$.
@@
This dataset consists of 47 stacks of 2D HASTE slices (3T, Siemens, TR=3-3.5s, TE=100ms, $1.25\!\times\!1.25\!\times\!3$$m$, FA=90) interleaved with 3D EPI navigator volumes (3T, Siemens, TR=25-46ms, 4-6mm, TE=12-22ms, flip angle=5) in 9 pregnant patients. Fetus GA ranges from 26 to 36 weeks, with mean GA $30.6\pm 2.9$ weeks. The average FOV size is $327\!\times\!327\!\times\!125$$m$.
Recall from Appendix A that slices are prescribed according to $\tilde{P}_{k}=T_{k}P_{k}$. To manually annotate GT poses in navigator volumes, we first algorithmically obtain $P_{k}$ values for all slices, using optimization-based slice-to-volume registration [xu2023]. Since all $\tilde{P}_{k}$ values are available to us as part of the acquisition parameters, we then compute $T_{k}=\tilde{P}_{k}P_{k}^{-1}$ as the GT pose for every navigator volume preceding the $k$th slice in the stack. Lastly, we manually correct the poses in each navigator volume.
Table 3: Performance statistics for translation estimation. Mean $\pm$ standard deviation for translation error ($m$) are reported. Best score is in bold. * indicates statistical significance compared to E(3)-Pose (p<0.05p\!<\!0.05, pairwise Wilcoxon test). We compare E(3)-Pose only to methods that estimate translation differently from E(3)-Pose. We evaluate the instance of E(3)-Pose trained on Research-Fetal since translation estimation relies only on brain segmentation. FireANTs runtime is 1.8s, prohibiting real-time deployment.
Research-Fetal test dHCP test Clinical-Young Navigators
FireANTs [jena2024]   2.0 $\pm$ 2.3* 0.5 $\mathbf{\pm}$ 1.6* 0.6 $\mathbf{\pm}$ 0.5* 4.2 $\pm$ 3.1*
EquiTrack [billot2024]   1.3 $\pm$ 0.9* 1.1 $\pm$ 0.7* 1.6 $\pm$ 1.6* 4.4 $\pm$ 3.2*
@@
Appendix F Baseline methods
We describe our implementations of the baseline methods and the modifications we make in order to optimize performance of baseline algorithms for our experiments.
FireANTs [jena2024]

### v5

@@ Appendix G Datasets
This dataset consists of whole-uterus EPI volumes in 153 pregnant volunteers (3T, Siemens, 3mm isotropic resolution, TR=2.9-4s, TE=32-47 ms, flip angle=90). Fetus GA ranges from 18 to 38 weeks, with mean GA $29.0\pm 5.0$ weeks. The average FOV size is $300\!\times\!300\!\times\!156$$m$.
@@
This dataset consists of fetal brain EPI volumes in 245 pregnant volunteers (3T, Philips Achieva, 2.2mm isotropic resolution, TR=2.2s, TE=60ms, flip angle=90) [37]. Fetus GA ranges from 20 to 38 weeks, with mean GA $28.9\pm 3.8$ weeks. The average FOV size is $143\!\times\!151\!\times\!108$mm. Volumes are cropped around the fetal head and are not whole-uterus volumes. Our manually annotated brain/eyes segmentations and poses for the dHCP dataset are publicly available on the project website for full transparency and future use.
@@
This dataset consists of routine clinical whole-uterus EPI volumes in 60 pregnant patients (3T, Siemens, 1.76-3.5mm voxels, TR=2.2-5s, TE=37ms, flip angle=90). There are 10 subjects for each GA in the range of 18 to 23 weeks. The average FOV size is $301\!\times\!301\!\times\!129$mm. Additionally, we note that while Clinical-Young is not representative of our target application of automated slice prescription, it is representative of routine clinical data acquired for retrospective head pose tracking, which promises to enable early disease detection. Specifically, there is well-supported evidence that fetal motion patterns correlate with neurodevelopmental disorders and intrauterine complications [47, 1].
@@
This dataset consists of 47 stacks of 2D HASTE slices (3T, Siemens, TR=3-3.5s, TE=100ms, $1.25\!\times\!1.25\!\times\!3$$m$, FA=90) interleaved with 3D EPI navigator volumes (3T, Siemens, TR=25-46ms, 4-6mm, TE=12-22ms, flip angle=5) in 9 pregnant volunteers. Fetus GA ranges from 26 to 36 weeks, with mean GA $30.6\pm 2.9$ weeks. The average FOV size is $327\!\times\!327\!\times\!125$mm. This dataset is most representative of automated slice prescription, and navigator volumes contain real spin history artifacts.
Figure 9: Segmentation study. Our segmentation network (trained on Research-Fetal) predicts accurate brain masks on all datasets, including challenging, out-of-distribution navigator volumes (left). Rotation estimation with the E(3)-CNN remains robust to larger segmentation errors observed in Navigators (top right), while translation estimation displays sensitivity to error (bottom right).
Table 4: Subject-level baseline comparisons for rotation error () in Navigators. Mean $\pm$ 95% confidence interval statistics are displayed. Subjects are denoted by gestational age (GA, weeks). Best score is in bold. * indicates statistical significance compared to E(3)-Pose (p<0.05p\!<\!0.05, permutation test). See Table 1 for additional details.
Subject (GA) S1 (28w) S2 (28w) S3 (32w) S4 (34w) S5 (36w) S6 (28w) S7 (26w) S8 (31w) S9 (28w)
FireANTs [36] 65.7 $\pm$ 8.2* 62.8 $\pm$ 8.6* 49.4 $\pm$ 12.9* 10.0 $\pm$ 1.2* 53.5 $\pm$ 8.1* 75.8 $\pm$ 14.3* 28.7 $\pm$ 6.4* 8.6 $\pm$ 1.7 24.0 $\pm$ 8.6*
Fetal-Align [33] 13.2 $\pm$ 4.1* 90.3 $\pm$ 5.3* 52.3 $\pm$ 10.5* 35.7 $\pm$ 8.6* 115.3 $\pm$ 8.4* 62.3 $\pm$ 11.2* 21.3 $\pm$ 4.2* 10.7 $\pm$ 3.9 29.4 $\pm$ 9.3*
Trained on Research-Fetal
EquiTrack [3] 40.1 $\pm$ 6.0* 63.9 $\pm$ 6.6* 37.7 $\pm$ 8.2* 45.5 $\pm$ 9.8* 45.3 $\pm$ 6.6* 97.5 $\pm$ 12.6* 24.2 $\pm$ 4.3* 28.5 $\pm$ 8.2* 21.3 $\pm$ 8.5*
3DPose-Net [60] 54.7 $\pm$ 6.0* 113.4 $\pm$ 5.6* 98.6 $\pm$ 8.7* 56.4 $\pm$ 6.7* 30.2 $\pm$ 2.8* 101.5 $\pm$ 9.3* 65.9 $\pm$ 7.6* 35.2 $\pm$ 6.1* 37.8 $\pm$ 4.5*
6DRep [16] 23.1 $\pm$ 2.5* 95.0 $\pm$ 7.0* 24.4 $\pm$ 3.9* 36.2 $\pm$ 7.8* 11.7 $\pm$ 1.8 52.6 $\pm$ 9.7* 46.3 $\pm$ 8.7* 18.5 $\pm$ 5.1* 19.2 $\pm$ 1.6*
RbR [25] 20.7 $\pm$ 2.9* 43.3 $\pm$ 3.8* 20.6 $\pm$ 3.8* 10.7 $\pm$ 1.2* 10.4 $\mathbf{\pm}$ 1.2 51.8 $\pm$ 10.1* 26.0 $\pm$ 3.0* 11.0 $\pm$ 0.9* 9.3 $\pm$ 0.8*
E(3)-Pose (ours) 6.4 $\mathbf{\pm}$ 0.5 13.6 $\mathbf{\pm}$ 1.5 10.6 $\mathbf{\pm}$ 3.4 6.6 $\mathbf{\pm}$ 0.6 10.7 $\pm$ 0.8 13.4 $\mathbf{\pm}$ 1.9 9.2 $\mathbf{\pm}$ 0.8 8.4 $\mathbf{\pm}$ 1.0 7.4 $\mathbf{\pm}$ 0.6
Trained on dHCP
EquiTrack [3] 32.9 $\pm$ 5.8* 66.3 $\pm$ 6.8* 65.2 $\pm$ 12.2* 75.5 $\pm$ 9.9* 76.7 $\pm$ 8.5* 109.4 $\pm$ 10.4* 18.8 $\pm$ 3.8 68.5 $\pm$ 13.4* 36.4 $\pm$ 10.9
3DPose-Net [60] 138.8 $\pm$ 3.1* 100.2 $\pm$ 3.5* 139.8 $\pm$ 4.7* 115.7 $\pm$ 4.7* 16.8 $\pm$ 1.7* 69.4 $\pm$ 4.7* 79.6 $\pm$ 3.0* 54.1 $\pm$ 3.7* 102.9 $\pm$ 2.9*
6DRep [16] 55.2 $\pm$ 4.2* 153.1 $\pm$ 4.2* 72.6 $\pm$ 2.9* 82.2 $\pm$ 12.7* 17.0 $\pm$ 1.1* 38.6 $\pm$ 3.3* 137.8 $\pm$ 3.6* 79.5 $\pm$ 4.2* 115.4 $\pm$ 7.7*
RbR [25] 45.3 $\pm$ 3.2* 121.0 $\pm$ 4.1* 100.8 $\pm$ 5.4* 41.1 $\pm$ 6.4* 23.1 $\pm$ 4.1* 118.6 $\pm$ 8.0* 145.6 $\pm$ 4.5* 86.3 $\pm$ 8.0* 37.1 $\pm$ 4.9*
E(3)-Pose (ours) 12.7 $\mathbf{\pm}$ 1.6 24.4 $\mathbf{\pm}$ 2.6 14.4 $\mathbf{\pm}$ 3.3 6.5 $\mathbf{\pm}$ 0.5 9.2 $\mathbf{\pm}$ 0.9 14.9 $\mathbf{\pm}$ 1.7 16.9 $\mathbf{\pm}$ 2.2 7.0 $\mathbf{\pm}$ 0.8 14.2 $\mathbf{\pm}$ 1.7
Table 5: Subject-level baseline comparisons for average absolute distance (AAD, $m$) in Navigators. Mean $\pm$ 95% confidence interval statistics are displayed. See Table 4 for description.
Subject (GA) S1 (28w) S2 (28w) S3 (32w) S4 (34w) S5 (36w) S6 (28w) S7 (26w) S8 (31w) S9 (28w)
FireANTs [36] 24.4 $\pm$ 2.5* 22.5 $\pm$ 2.6* 22.5 $\pm$ 4.7* 6.2 $\pm$ 0.7* 27.5 $\pm$ 3.8* 33.1 $\pm$ 4.8* 12.1 $\pm$ 2.1* 5.8 $\pm$ 0.9 11.2 $\pm$ 3.0*
Fetal-Align [33] 7.4 $\pm$ 1.4* 36.2 $\pm$ 2.0* 27.4 $\pm$ 4.8* 18.8 $\pm$ 3.8* 49.5 $\pm$ 3.3* 29.0 $\pm$ 4.5* 9.5 $\pm$ 1.6* 6.0 $\pm$ 1.6 13.4 $\pm$ 3.8*
Trained on Research-Fetal
EquiTrack [3] 17.5 $\pm$ 2.1* 25.7 $\pm$ 2.2* 19.9 $\pm$ 3.2* 21.9 $\pm$ 3.8* 23.7 $\pm$ 2.8* 41.4 $\pm$ 4.1* 11.4 $\pm$ 1.5* 13.0 $\pm$ 2.9* 9.3 $\pm$ 2.9*
3DPose-Net [60] 23.1 $\pm$ 2.1* 42.1 $\pm$ 1.6* 43.2 $\pm$ 2.6* 29.7 $\pm$ 2.6* 17.5 $\pm$ 1.4* 43.1 $\pm$ 2.7* 24.3 $\pm$ 2.3* 16.3 $\pm$ 2.2* 17.2 $\pm$ 1.9*
6DRep [16] 11.4 $\pm$ 1.1* 36.5 $\pm$ 2.2* 14.2 $\pm$ 2.0* 18.7 $\pm$ 3.2* 7.9 $\pm$ 0.9 25.9 $\pm$ 3.5* 16.5 $\pm$ 2.5* 9.0 $\pm$ 2.0* 9.1 $\pm$ 0.7*
RbR [25] 10.4 $\pm$ 1.1* 19.4 $\pm$ 1.4* 12.2 $\pm$ 1.8* 7.4 $\pm$ 0.6* 7.3 $\mathbf{\pm}$ 0.7 25.8 $\pm$ 4.1* 11.0 $\pm$ 1.1* 6.3 $\pm$ 0.5* 5.0 $\pm$ 0.4*
E(3)-Pose (ours) 5.0 $\mathbf{\pm}$ 0.3 7.6 $\mathbf{\pm}$ 0.7 7.5 $\mathbf{\pm}$ 1.7 5.5 $\mathbf{\pm}$ 0.2 7.4 $\pm$ 0.5 10.1 $\mathbf{\pm}$ 1.2 4.8 $\mathbf{\pm}$ 0.3 5.1 $\mathbf{\pm}$ 0.5 4.1 $\mathbf{\pm}$ 0.3
Trained on dHCP
EquiTrack [3] 14.8 $\pm$ 2.2* 26.1 $\pm$ 2.3* 29.8 $\pm$ 4.7* 36.2 $\pm$ 4.1* 36.8 $\pm$ 3.4* 45.9 $\pm$ 2.7* 8.3 $\pm$ 1.3 25.2 $\pm$ 4.4* 14.7 $\pm$ 3.7
3DPose-Net [60] 49.5 $\pm$ 0.6* 38.2 $\pm$ 1.1* 60.1 $\pm$ 1.0* 54.0 $\pm$ 1.5* 10.5 $\pm$ 0.9* 33.8 $\pm$ 1.8* 31.2 $\pm$ 1.0* 26.5 $\pm$ 1.5* 40.9 $\pm$ 0.9*
6DRep [16] 23.9 $\pm$ 1.6* 49.8 $\pm$ 1.0* 36.0 $\pm$ 1.3* 35.2 $\pm$ 4.7* 10.8 $\pm$ 10.6* 20.4 $\pm$ 1.5* 44.5 $\pm$ 0.7* 36.8 $\pm$ 1.7* 45.0 $\pm$ 2.0*
RbR [25] 20.4 $\pm$ 1.2* 43.5 $\pm$ 0.9* 48.3 $\pm$ 1.8* 21.7 $\pm$ 2.8* 13.2 $\pm$ 1.8* 48.5 $\pm$ 2.4* 45.4 $\pm$ 0.8* 34.7 $\pm$ 2.7* 17.1 $\pm$ 1.8*
E(3)-Pose (ours) 7.3 $\mathbf{\pm}$ 0.7 11.9 $\mathbf{\pm}$ 1.0 9.2 $\mathbf{\pm}$ 1.6 5.4 $\mathbf{\pm}$ 0.3 6.7 $\mathbf{\pm}$ 0.6 10.6 $\mathbf{\pm}$ 1.1 7.9 $\mathbf{\pm}$ 0.9 4.5 $\mathbf{\pm}$ 0.4 7.2 $\mathbf{\pm}$ 0.8
Table 6: Subject-level baseline comparisons for translation error ($m$) in Navigators. Mean $\pm$ 95% confidence interval statistics are displayed. See Table 4 for description.
Subject (GA) S1 (28w) S2 (28w) S3 (32w) S4 (34w) S5 (36w) S6 (28w) S7 (26w) S8 (31w) S9 (28w)
FireANTs [36] 3.6 $\mathbf{\pm}$ 0.3* 4.6 $\pm$ 0.4* 5.8 $\pm$ 1.1* 2.7 $\mathbf{\pm}$ 0.3* 4.6 $\pm$ 0.4* 8.1 $\pm$ 1.0* 3.0 $\pm$ 0.2* 3.4 $\pm$ 0.3* 3.5 $\pm$ 0.4*
EquiTrack [3] 4.0 $\pm$ 0.3 5.6 $\pm$ 0.5* 5.1 $\pm$ 1.0 3.7 $\pm$ 0.5 3.9 $\pm$ 0.4* 7.5 $\pm$ 1.0 4.5 $\pm$ 0.3* 3.1 $\pm$ 0.3* 2.1 $\pm$ 0.3*
E(3)-Pose (ours) 4.0 $\pm$ 0.3 4.2 $\mathbf{\pm}$ 0.4 5.0 $\mathbf{\pm}$ 1.0 3.9 $\pm$ 0.2 3.6 $\mathbf{\pm}$ 0.4 7.4 $\mathbf{\pm}$ 0.9 2.7 $\mathbf{\pm}$ 0.2 2.6 $\mathbf{\pm}$ 0.2 2.0 $\mathbf{\pm}$ 0.3
Table 7: Subject-level ablation study for rotation error () in Navigators. Mean $\pm$ 95% confidence interval statistics are displayed. See Table 4 for description.
Subject (GA) S1 (28w) S2 (28w) S3 (32w) S4 (34w) S5 (36w) S6 (28w) S7 (26w) S8 (31w) S9 (28w)
Trained on Research-Fetal
E(3)-Pose (ours) 6.4 $\mathbf{\pm}$ 0.5 13.6 $\mathbf{\pm}$ 1.5 10.6 $\mathbf{\pm}$ 3.4 6.6 $\pm$ 0.6 10.7 $\pm$ 0.8 13.4 $\mathbf{\pm}$ 1.9 9.2 $\pm$ 0.8 8.4 $\pm$ 1.0 7.4 $\pm$ 0.6
Standard CNN 10.8 $\pm$ 0.4* 32.4 $\pm$ 4.0* 19.8 $\pm$ 4.1* 12.7 $\pm$ 2.3* 10.2 $\pm$ 0.9 22.5 $\pm$ 4.3* 29.0 $\pm$ 3.7* 11.7 $\pm$ 1.7* 7.5 $\pm$ 0.6
no pseudovector 9.4 $\pm$ 0.9* 15.8 $\pm$ 2.4* 11.2 $\pm$ 3.2 5.8 $\mathbf{\pm}$ 0.5* 12.6 $\pm$ 1.0* 18.0 $\pm$ 4.0* 8.7 $\mathbf{\pm}$ 1.0 6.2 $\mathbf{\pm}$ 0.6* 8.6 $\pm$ 0.6*
$h(R)\!=\!e_{y}\!\oplus\!e_{z}$ 7.5 $\pm$ 0.6* 19.7 $\pm$ 2.9* 11.0 $\pm$ 3.6* 6.5 $\pm$ 0.8 9.6 $\mathbf{\pm}$ 0.7* 20.8 $\pm$ 2.5* 14.0 $\pm$ 2.2* 9.7 $\pm$ 0.8 7.3 $\mathbf{\pm}$ 0.7
$\mathcal{L}_{\phi}=\!\sum_{k\in\{x,y,z\}}|\sin{(\theta_{k}/2)}|$ 7.9 $\pm$ 0.5* 19.2 $\pm$ 2.2* 12.8 $\pm$ 3.6* 7.4 $\pm$ 0.9 11.1 $\pm$ 0.9 19.8 $\pm$ 3.6* 10.9 $\pm$ 1.3* 10.5 $\pm$ 0.8* 8.7 $\pm$ 0.6*
geodesic loss 7.2 $\pm$ 0.5* 25.8 $\pm$ 4.6* 12.6 $\pm$ 3.4* 6.4 $\pm$ 0.5 12.9 $\mathbf{\pm}$ 1.6* 23.1 $\pm$ 5.9* 14.1 $\pm$ 3.7* 6.2 $\mathbf{\pm}$ 0.7* 8.7 $\pm$ 1.1*
no artifact augmentation 11.6 $\pm$ 1.5* 22.7 $\pm$ 3.1* 11.8 $\pm$ 3.6 15.5 $\pm$ 2.5* 13.0 $\pm$ 1.0* 35.8 $\pm$ 7.0* 15.5 $\pm$ 3.2* 11.9 $\pm$ 2.2* 12.3 $\pm$ 4.2*
Trained on dHCP
E(3)-Pose (ours) 12.7 $\mathbf{\pm}$ 1.6 24.4 $\mathbf{\pm}$ 2.6 14.4 $\mathbf{\pm}$ 3.3 6.5 $\mathbf{\pm}$ 0.5 9.2 $\pm$ 0.9 14.9 $\mathbf{\pm}$ 1.7 16.9 $\mathbf{\pm}$ 2.2 7.0 $\pm$ 0.8 14.2 $\mathbf{\pm}$ 1.7
Standard CNN 93.9 $\pm$ 6.9* 108.0 $\pm$ 3.7* 147.0 $\pm$ 4.7* 28.4 $\pm$ 3.9* 11.7 $\pm$ 1.0* 105.4 $\pm$ 2.9* 111.8 $\pm$ 6.0* 80.6 $\pm$ 12.1* 86.3 $\pm$ 4.1*
no pseudovector 13.1 $\pm$ 1.9 45.3 $\pm$ 5.3* 21.9 $\pm$ 4.4* 8.9 $\pm$ 1.1* 9.4 $\pm$ 0.9 32.7 $\pm$ 9.6* 21.9 $\pm$ 3.2* 5.9 $\mathbf{\pm}$ 0.7* 23.7 $\pm$ 3.6*
$h(R)\!=\!e_{y}\!\oplus\!e_{z}$ 11.2 $\pm$ 1.2 43.7 $\pm$ 6.1* 15.0 $\pm$ 5.0 8.8 $\pm$ 0.9* 11.4 $\pm$ 1.5* 38.1 $\pm$ 8.7* 66.1 $\pm$ 8.3* 10.7 $\pm$ 3.1* 25.3 $\pm$ 5.0*
$\mathcal{L}_{\phi}=\!\sum_{k\in\{x,y,z\}}|\sin{(\theta_{k}/2)}|$ 9.4 $\pm$ 0.9* 28.0 $\pm$ 4.3 16.7 $\pm$ 5.2 7.7 $\pm$ 0.7* 8.7 $\mathbf{\pm}$ 0.8 29.3 $\pm$ 8.4* 42.5 $\pm$ 7.7* 9.5 $\pm$ 3.1 15.9 $\pm$ 1.5
geodesic loss 18.9 $\pm$ 4.8* 47.8 $\pm$ 7.1* 47.5 $\pm$ 14.5* 11.1 $\pm$ 0.8* 13.8 $\pm$ 1.1* 35.5 $\pm$ 8.8* 28.6 $\pm$ 6.7* 32.3 $\pm$ 11.1* 22.5 $\pm$ 7.1*
no artifact augmentation 40.9 $\pm$ 6.7* 47.6 $\pm$ 5.8* 34.3 $\pm$ 8.3* 22.5 $\pm$ 6.1* 16.5 $\pm$ 2.2* 55.0 $\pm$ 9.7* 26.9 $\pm$ 5.0* 19.5 $\pm$ 4.4* 28.2 $\pm$ 4.7*
Table 8: Subject-level ablation study for average absolute distance (AAD, $m$) in Navigators. Mean $\pm$ 95% confidence interval statistics are displayed. See Table 4 for description.
Subject (GA) S1 (28w) S2 (28w) S3 (32w) S4 (34w) S5 (36w) S6 (28w) S7 (26w) S8 (31w) S9 (28w)
Trained on Research-Fetal
E(3)-Pose (ours) 5.0 $\mathbf{\pm}$ 0.3 7.6 $\mathbf{\pm}$ 0.7 7.5 $\mathbf{\pm}$ 1.7 5.5 $\pm$ 0.2 7.4 $\pm$ 0.5 10.1 $\mathbf{\pm}$ 1.2 4.8 $\pm$ 0.3 5.1 $\pm$ 0.5 4.1 $\mathbf{\pm}$ 0.3
Standard CNN 6.3 $\pm$ 0.3* 14.9 $\pm$ 1.5* 12.0 $\pm$ 1.9* 8.9 $\pm$ 1.1* 7.6 $\pm$ 0.5 14.0 $\pm$ 1.9* 11.9 $\pm$ 1.2* 6.5 $\pm$ 0.8* 4.3 $\pm$ 0.4
no pseudovector 5.9 $\pm$ 0.4* 8.7 $\pm$ 0.9* 8.5 $\pm$ 1.5 5.3 $\mathbf{\pm}$ 0.2 8.7 $\pm$ 0.6* 11.9 $\pm$ 1.8* 4.7 $\mathbf{\pm}$ 0.4 4.1 $\mathbf{\pm}$ 0.3* 4.8 $\pm$ 0.3*
$h(R)\!=\!e_{y}\!\oplus\!e_{z}$ 5.5 $\pm$ 0.3* 10.1 $\pm$ 1.1* 8.2 $\pm$ 1.6 5.7 $\pm$ 0.4 7.3 $\mathbf{\pm}$ 0.5 13.2 $\pm$ 1.3* 6.5 $\pm$ 0.7* 5.5 $\pm$ 0.4 4.1 $\mathbf{\pm}$ 0.3
$\mathcal{L}_{\phi}=\!\sum_{k\in\{x,y,z\}}|\sin{(\theta_{k}/2)}|$ 5.4 $\pm$ 0.3* 9.7 $\pm$ 0.9* 8.4 $\pm$ 1.7* 5.8 $\pm$ 0.5 7.6 $\pm$ 0.6 12.7 $\pm$ 1.6* 5.5 $\pm$ 0.5* 5.9 $\pm$ 0.4* 4.8 $\pm$ 0.3*
geodesic loss 5.1 $\pm$ 0.3 11.7 $\pm$ 1.6* 8.4 $\pm$ 1.6* 5.3 $\mathbf{\pm}$ 0.3 8.5 $\pm$ 10.8* 14.1 $\pm$ 2.4* 6.4 $\pm$ 1.2* 4.1 $\mathbf{\pm}$ 0.4* 4.7 $\pm$ 0.5*
no artifact augmentation 6.8 $\pm$ 0.6* 11.2 $\pm$ 1.2* 8.7 $\pm$ 1.6* 9.5 $\pm$ 21.1* 8.9 $\pm$ 0.6* 18.9 $\pm$ 2.6* 7.1 $\pm$ 1.1* 6.6 $\pm$ 1.0* 6.1 $\pm$ 1.6*
Trained on dHCP
E(3)-Pose (ours) 7.3 $\pm$ 0.7 11.9 $\mathbf{\pm}$ 1.0 9.2 $\mathbf{\pm}$ 1.6 5.4 $\mathbf{\pm}$ 0.3 6.7 $\pm$ 0.6 10.6 $\mathbf{\pm}$ 1.1 7.9 $\mathbf{\pm}$ 0.9 4.5 $\pm$ 0.4 7.2 $\mathbf{\pm}$ 0.8
Standard CNN 35.0 $\pm$ 2.0* 39.4 $\pm$ 0.8* 57.0 $\pm$ 0.7* 16.1 $\pm$ 1.9* 8.0 $\pm$ 0.6* 47.1 $\pm$ 1.1* 38.6 $\pm$ 1.2* 32.0 $\pm$ 4.0* 39.3 $\pm$ 1.6*
no pseudovector 7.3 $\pm$ 0.8 19.9 $\pm$ 2.0* 13.0 $\pm$ 2.0* 6.5 $\pm$ 0.6* 6.8 $\pm$ 0.6 16.9 $\pm$ 3.2* 9.6 $\pm$ 1.2* 4.0 $\mathbf{\pm}$ 0.3* 11.5 $\pm$ 1.5*
$h(R)\!=\!e_{y}\!\oplus\!e_{z}$ 6.8 $\pm$ 0.5 18.2 $\pm$ 2.0* 9.4 $\pm$ 2.1 6.4 $\pm$ 0.4* 7.9 $\pm$ 0.9* 20.3 $\pm$ 3.5* 24.1 $\pm$ 2.5* 6.0 $\pm$ 1.1* 12.2 $\pm$ 2.1*
$\mathcal{L}_{\phi}=\!\sum_{k\in\{x,y,z\}}|\sin{(\theta_{k}/2)}|$ 6.1 $\mathbf{\pm}$ 0.4* 12.7 $\pm$ 1.5 10.1 $\pm$ 2.3 6.0 $\pm$ 0.3* 6.4 $\mathbf{\pm}$ 0.5 16.0 $\pm$ 3.1* 16.1 $\pm$ 2.4* 5.5 $\pm$ 1.1* 8.2 $\pm$ 0.7*
geodesic loss 9.2 $\pm$ 1.5* 19.0 $\pm$ 2.3* 20.7 $\pm$ 5.3* 7.2 $\pm$ 0.4* 8.9 $\pm$ 0.6* 18.7 $\pm$ 3.2* 11.0 $\pm$ 1.9* 12.4 $\pm$ 3.4* 10.2 $\pm$ 2.4*
no artifact augmentation 17.4 $\pm$ 2.3* 20.2 $\pm$ 2.0* 18.3 $\pm$ 3.6* 12.5 $\pm$ 2.5* 10.5 $\pm$ 1.1* 26.5 $\pm$ 3.8* 11.3 $\pm$ 1.8* 9.7 $\pm$ 1.8* 13.6 $\pm$ 2.0*
@@
Appendix H Baseline methods
We describe our implementations of the baseline methods and the modifications we make in order to optimize performance of baseline algorithms for our experiments. All methods that rely on network training were trained on a Nvidia RTX 6000 Ada GPU with batch size of either 4 or 5, unless otherwise stated.
FireANTs [36]